# Lightweight Dataset Pruning without Full Training via Example Difficulty and Prediction Uncertainty

### v2

@@ B.4 Effectiveness of Beta Sampling
We study the impact of our Beta sampling on existing score metrics. We apply our Beta sampling strategy to forgetting, EL2N, and Dyn-Unc scores of CIFAR10 and 100. By comparing Beta sampling with the vanilla threshold pruning using scores, we observe that prior score-based methods become competitive, outperforming random pruning when Beta sampling is adjusted.
@@
We also study the impact of our pruning strategy with DUAL score combined with Beta sampling. We compare different sampling strategies $i.e.$ vanilla thresholding, stratified sampling (Zheng et al., 2022), and our proposed Beta sampling on CIFAR10 and 100, at 80% and 90% pruning rates. We observe that our proposed method mostly performs the best, especially with the high pruning ratio.
@@ C.1 Visualization of Selected Data with Beta Sampling
Here we illustrate the sampling probability of being selected into coreset, selected samples, and pruned samples in each figure when using the DUAL score combined with Beta sampling. As the pruning ratio increases, we focus on including easier samples.
(a) CIFAR-100 at pruning ratio 30%
(b) CIFAR-100 at pruning ratio 50%
(c) CIFAR-100 at pruning ratio 70%
(d) CIFAR-100 at pruning ratio 90%
@@ C.2 Algorithm of Proposed Pruning Method
        ¯$\bar{\mathbb{P}}_{k}({\bm{x}}_{i},y_{i})\leftarrow\frac{1}{J}\sum_{j=0}^{J-1}\mathbb{P}_{k+j}(y_{i}\mid{\bm{x}}_{i})$
        $\mathbb{U}_{k}({\bm{x}}_{i},y_{i})\leftarrow\sqrt{\frac{1}{J-1}\sum_{j=0}^{J-1}\left[\mathbb{P}_{k+j}(y_{i}\mid{\bm{x}}_{i})-\bar{\mathbb{P}}_{k}({\bm{x}}_{i},y_{i})\right]^{2}}$
@@
Algorithm 1 DUAL pruning + $\beta$-sampling
@@ Appendix D Theoretical Results
$V_{t;J}^{(1)}\coloneq\text{Var}\left[\left\{\sigma\left(y_{\tau}^{(1)}\right)\right\}_{\tau=t}^{t+J-1}\right]$ $=\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(1)}\right)-\sigma\left(y_{t+k}^{(1)}\right)\right]^{2}$
$>\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(2)}\right)-\sigma\left(y_{t+k}^{(2)}\right)\right]^{2}$
$=\text{Var}\left[\left\{\sigma\left(y_{\tau}^{(2)}\right)\right\}_{\tau=t}^{t+J-1}\right]\eqcolon V_{t;J}^{(2)}.$
@@
$V_{t;J}^{(1)}\left(1-\mu_{t;J}^{(1)}\right)$
$=\left[\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(1)}\right)-\sigma\left(y_{t+k}^{(1)}\right)\right]^{2}\right]\left[1-\frac{1}{J}\sum_{\tau=t}^{t+J-1}\sigma\left(y_{\tau}^{(1)}\right)\right]$
$>\left[\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(2)}\right)-\sigma\left(y_{t+k}^{(2)}\right)\right]^{2}\right]\left[1-\frac{1}{J}\sum_{\tau=t}^{t+J-1}\sigma\left(y_{\tau}^{(2)}\right)\right]$
$=V_{t;J}^{(2)}\left(1-\mu_{t;J}^{(2)}\right).$
@@
$\begin{split}&V_{t;J}^{(1)}\left(1-\mu_{t;J}^{(1)}\right)\\ &=\left[\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(1)}\right)-\sigma\left(y_{t+k}^{(1)}\right)\right]^{2}\right]\left[\frac{1}{J}\sum_{\tau=t}^{t+J-1}1-\sigma\left(y_{\tau}^{(1)}\right)\right]\\ &>\left[\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(2)}\right)-\sigma\left(y_{t+k}^{(2)}\right)\right]^{2}\right]\left[\frac{1}{J}\sum_{\tau=t}^{t+J-1}1-\sigma\left(y_{\tau}^{(2)}\right)\right]\\ &=V_{t;J}^{(2)}\left(1-\mu_{t;J}^{(2)}\right).\end{split}$ (13)
@@
$\left[\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(1)}\right)-\sigma\left(y_{t+k}^{(1)}\right)\right]^{2}\right]\left[\frac{1}{J}\sum_{\tau=t}^{t+J-1}1-\sigma\left(y_{\tau}^{(1)}\right)\right]$ (14)
$>\left[\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(2)}\right)-\sigma\left(y_{t+k}^{(2)}\right)\right]^{2}\right]\left[\frac{1}{J}\sum_{\tau=t}^{t+J-1}1-\sigma\left(y_{\tau}^{(2)}\right)\right].$

### v3

@@ B.6 Ablation Study on Beta Sampling
We study the impact of our Beta sampling on existing score metrics. We apply our Beta sampling strategy to forgetting, EL2N, and Dyn-Unc scores of CIFAR10 and 100. By comparing Beta sampling with the vanilla threshold pruning using scores, with results presented in Tables 21 and 22, we observe that prior score-based methods become competitive, outperforming random pruning when Beta sampling is adjusted.
@@
We also study the impact of our pruning strategy with DUAL score combined with Beta sampling. We compare different sampling strategies: vanilla thresholding, stratified sampling (Zheng et al., 2022), and Beta sampling on CIFAR10 and 100, at 80% and 90% pruning rates. The results, presented in Table 23 indicate that our proposed Beta sampling mostly performs the best, especially with the high pruning ratio.
@@ C.1 Visualization of Selected Data with Beta Sampling
We provide visualizations of Beta sampling in Figure 21. This figure illustrates, in respective columns: (i) sample selection probabilities for the coreset; (ii) examples of selected samples when forming a 70%, 30% and 10% coreset; and (iii) the samples pruned at these corresponding rates. The visualization shows that Beta sampling increasingly favors the selection of easier samples for the coreset as the pruning ratio increases.
(a) CIFAR-100 with pruning ratio 30%
(b) CIFAR-100 with pruning ratio 70%
(c) CIFAR-100 with pruning ratio 90%
@@ C.2 Algorithm of Proposed Pruning Method
        ¯$\bar{\mathbb{P}}_{k}({\bm{x}}_{i},y_{i})\leftarrow\frac{1}{J}\sum_{j=0}^{J-1}\mathbb{P}_{k+j}(y_{i}\mid{\bm{x}}_{i})$Example Difficulty}
        $\mathbb{U}_{k}({\bm{x}}_{i},y_{i})\leftarrow\sqrt{\frac{1}{J-1}\sum_{j=0}^{J-1}\left[\mathbb{P}_{k+j}(y_{i}\mid{\bm{x}}_{i})-\bar{\mathbb{P}}_{k}({\bm{x}}_{i},y_{i})\right]^{2}}$Prediction Uncertainty}
@@
Algorithm 1 DUAL pruning + $\beta$ sampling
@@ Appendix D Theoretical Results
$S_{t;J}^{(1)}\coloneq\sqrt{\text{Var}\left[\left\{\sigma\left(y_{\tau}^{(1)}\right)\right\}_{\tau=t}^{t+J-1}\right]}$ $=\sqrt{\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(1)}\right)-\sigma\left(y_{t+k}^{(1)}\right)\right]^{2}}$
$>\sqrt{\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(2)}\right)-\sigma\left(y_{t+k}^{(2)}\right)\right]^{2}}$
$=\sqrt{\text{Var}\left[\left\{\sigma\left(y_{\tau}^{(2)}\right)\right\}_{\tau=t}^{t+J-1}\right]}\eqcolon S_{t;J}^{(2)}.$
@@
$S_{t;J}^{(1)}\left(1-\mu_{t;J}^{(1)}\right)$
$=\sqrt{\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(1)}\right)-\sigma\left(y_{t+k}^{(1)}\right)\right]^{2}}\left[1-\frac{1}{J}\sum_{\tau=t}^{t+J-1}\sigma\left(y_{\tau}^{(1)}\right)\right]$
$>\sqrt{\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(2)}\right)-\sigma\left(y_{t+k}^{(2)}\right)\right]^{2}}\left[1-\frac{1}{J}\sum_{\tau=t}^{t+J-1}\sigma\left(y_{\tau}^{(2)}\right)\right]$
$=S_{t;J}^{(2)}\left(1-\mu_{t;J}^{(2)}\right).$
@@
$S_{t;J}^{(1)}\left(1-\mu_{t;J}^{(1)}\right)$ (13)
$=\sqrt{\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(1)}\right)-\sigma\left(y_{t+k}^{(1)}\right)\right]^{2}}\left[\frac{1}{J}\sum_{\tau=t}^{t+J-1}1-\sigma\left(y_{\tau}^{(1)}\right)\right]$ (14)
$>\sqrt{\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(2)}\right)-\sigma\left(y_{t+k}^{(2)}\right)\right]^{2}}\left[\frac{1}{J}\sum_{\tau=t}^{t+J-1}1-\sigma\left(y_{\tau}^{(2)}\right)\right]$ (15)
$=S_{t;J}^{(2)}\left(1-\mu_{t;J}^{(2)}\right).$ (16)
@@
$\sqrt{\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(1)}\right)-\sigma\left(y_{t+k}^{(1)}\right)\right]^{2}}\left[\frac{1}{J}\sum_{\tau=t}^{t+J-1}1-\sigma\left(y_{\tau}^{(1)}\right)\right]$ (17)
$>\sqrt{\frac{1}{J(J-1)}\sum_{k=0}^{J-2}\sum_{l=k+1}^{J-1}\left[\sigma\left(y_{t+l}^{(2)}\right)-\sigma\left(y_{t+k}^{(2)}\right)\right]^{2}}\left[\frac{1}{J}\sum_{\tau=t}^{t+J-1}1-\sigma\left(y_{\tau}^{(2)}\right)\right]$